OpenStudy (anonymous):

I am stuck help!!!! 2-(4/x^2) =2/x 2=2/x+4x^2 I then got 2=2x/x^2 + 4/x^2 2= 2x+4/x^2 What to do next I dont know what to do with the x^2

OpenStudy (anonymous):

\[2-\frac{ 4 }{ x ^{2} }=\frac{ 2 }{ x }\] Is that the original problem?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

It would be easiest if you find a common denominator first. Then you can multiply each term by the common denominator and clear all the fractions.

OpenStudy (anonymous):

X^2 IS the common deniminator

OpenStudy (anonymous):

Correct. So multiply each of the terms by x^2, then that will clear your denominators.

OpenStudy (anonymous):

I don't know how to do that. I get what u mean but if how to set it up

OpenStudy (anonymous):

2 times x^2 = 2x^2 4/x^2 times x^2 will cancel the x^2 and just = 4 2/x times x^2 will cancel one x so = 2x So you'd have \[2x^{2}-4=2x\]

OpenStudy (anonymous):

Do I just factor now

OpenStudy (anonymous):

Get all terms on one side, then factor. Yep!

OpenStudy (anonymous):

OK so it would be 2 (x+1)(x-2)